5 years ago, Brian was twice as old as John.If Brian is 7 Years older than John, how old was Brian 9 years ago?
Answer by ankor@dixie-net.com(22740)

Let b = Brian's present age
Let j = John's present age
:
Write an equation for each statement:
;
"5 years ago, Brian was twice as old as John."
b - 5 = 2(j-5)
b - 5 = 2j - 10
b = 2j - 10 + 5
b = 2j - 5
;
"If Brian is 7 Years older than John,"
b = j + 7
Replace b with (2j-5)
2j - 5 = j + 7
2j - j = 7 + 5
j = 12 yr is John's present age
then
12 + 7 = 19 yrs is Brian's present age
:
how old was Brian 9 years ago?
19 - 9 = 10 yrs old, was Brian, 9 yrs ago
:
:
Check solutions for present age in the statement
"5 years ago, Brian was twice as old as John."
19 - 5 = 2(12-5)
14 = 2(7)